how would I show my work on this problem ?
What is the interest earned on $24,000 at 8 1/2 for 2 years ?

Answer by stanbon(75887)

What is the interest earned on $24,000 at 8 1/2% for 2 years ?
---
Simple Interest::
I = P*r*t
I = 24000*0.085*2 = $4080
-----------------------
Compound Interest::
I(t) = P(1 + (r/n))^(n^t) - P
I(2) = 24000(1+(0.085/1))^(1*2) - 24000
I(2) = 24000(1.085)^2-1) = 24000(0.17723) = $4253
--------------
Cheers,
Stan H.
-----------